What is the angle between the hour and the minute hands if the time is 9:20?
Answer by stanbon(75887)

What is the angle between the hour and the minute hands if the time is 9:20?
Draw the picture.
---
The smaller angle is (5/12)360 = 5*30 = 150 degrees
The larger angle is (7/12)360) = 7*30 = 210 degrees
=====================================================
Cheers,
Stan H.